OpenStudy (anonymous):

Integration help: Q: x^2/(x^6+64)dx

OpenStudy (ageta):

\[\int\limits_{}^{} Q:\frac{ x^2 }{ 6x+64 }dx\]

OpenStudy (ageta):

does that help

OpenStudy (anonymous):

thats the question

OpenStudy (zarkon):

\[\int \frac{ x^2 }{ x^6+64 }dx= \int \frac{ x^2 }{ (x^3)^2+64 }dx \]

OpenStudy (anonymous):

yep

OpenStudy (ageta):